OpenStudy (anonymous):

Solve for x --> x=7/8 divided by 1/4 . Express your answer in simplest form

jagr2713 (jagr2713):

can you send pic?

OpenStudy (crashonce):

7/8divided by 1/4=7/8 x 4 = 7/2

OpenStudy (midhun.madhu1987):